PIPVIC

Videoconference 20/03/98 10.00-11.00

Attendees:-
Julian Chesterfield, UCL (JC)
Louise Clark, UCL (LC)
David Hearnshaw, U Westminster (DH)
Paul Hogan, UCL (PH)
John Buckett, Exeter (JB)
Kris Hasler, UCL (KH)
Adriano Ventencelli, Aber (AV)
Angela Sasse, UCL (MAS)
Dave Price, Aber (DAP)
Henry Hughes, UKERNA (HRH)
 
 
Apologies:-
Roy Bennett, UCL (RB)
Colin Perkins, UCL (CSP)
 
Meeting Minutes:-
 

1. Connection and tool check


Exeter, Aber, Westminster(except vic) - fine
Essex cut off

2. Contract update


MAS: Sent bill for first 50% to UKERNA. Copy of contract has been sent to UKERNA. HRH has not received contract yet, UKERNA will pay once they have signed contract.

3. Update on structured activities


a) Teaching:
 
- -Spanish for Lawyers
Evaluation session takes place tomorrow - start 10.45 at UCL
JB: Exeter - confirmed.

- -Italian
AV: Last two sessions last night.
GS: Went well, no network problems. Tried using both
WB and NTE. PC used last night, and quality was very good.

DAP: new version of rat seems to work much better than previous versions. However, more testing needed. A few problems after a while, but not as bad as previously

Evaluation session 28th April. UCL to join in.

- -Polish - evaluation session set for 21st april
 
- -Essex-SSEES Social Science Teaching
LP: delay initially with slides, session went well after that.

JC: Problems with RAT freezing. Mediaboard needs to be set up in advance.
Went well eventually. Long delay - 20-25 minutes to transmit slides. Audio quality much better, but RAT froze on occassions 2-3 times. Crashed when adjusting volume control. Students very positive. Mediaboard wasn't used in the end - material in media board to be used for postponed session. Slides should be loaded half an hour before the start of session.

MAS: slides should be prepared well in advance of session and downloaded beforehand.
 

b) Seminars - upcoming events

Ian Campbell

Craig Rodine ain ?? OU
 
 

4. PC tools - update/review of new releases


LC: new release seems much improved.
 
KH: problem with rat - it hits 100 per cent CPU loading, and must be shut down and restarted.

ACTION (KH, JC) - check which cards work with new version of rat ASAP

UCL is still working hard on weeding out problems with rat. Scheduling and buffer problems solved so far.

KH: New version of vic on KH's homepage which is a vast improvement on previous version.
Please download this new version.:-
http://www.cs.ucl.ac.uk/staff/K.Hasler/shrimp/bin
It is now possible to enlarge windows on PCs without loading the CPU fully.

HH: can the latest versions of rat & vic be made available on the PIPVIC Web pages?

ACTION (LC, KH) pointer from PIPVIC webpage to latest version of PC tools - with dates and times...
 

5. AOB

MAS - No teaching sessions now for some weeks, please test tools extensively at all sites.

ACTION (All sites - Julian in charge) - Weekly sessions testing PC software SYSTEMATICALLY.

AV - About to buy new PCs, would like to know which hardware to buy as soon as we know. The sooner the better.

6. Outstanding Action points


ACTION 98/03/20-1: MAS to check up on contract situation.
Done

ACTION 98/03/20: PH & MAS to identify some possible dates in late April for
(language) tutor workshop and email to tutors.

ACTION 98/03/13-5 (KH): Look into interoperability of wb and wbd

ACTION 98/03/06-5: More specific info on how best to prepare slides for
the whiteboard. MAS/LC/KH to look at this and prepare guidelines for size, shape, fonts, etc. - in process, proposals to come

ACTION 98/03/06 - DH and MAS to do series on hands-on experiences of using
videoconferencing technology for teaching.
Dates: April 23,30 May 7, 14, 21, 28
 
ACTION 98/03/20 - DP to file report on packet loss problems in last weeks meeting. - done..
DAP: The nosc responded well, and were as puzzled as we
were. The fault "appeared" to have cleared and I had a phone call from
Freddie Williams at the NOSC yesterday (thursday) enquiring about
whether of not the fault had cleared. This was all handled under
job number JOD1226. I said "clear as far as we can see", but
our big meeting was this morning. We also note that the NOSC
are now very actively sending out TICKETS for faults/changes
affecting the MBONE. A Ticket arrived last night reporting
a hardware problem at Telehouse where this fault appears still to be.